5g of acetylsalicylic acid (aspirin) was obtained from 2.4g of salicylic acid. Calculate % yield, if the theoretical yield is 3.6g.

1
Expert's answer
2022-05-13T10:33:03-0400

Aspirin is prepared by the reaction between acetic anhydride and salicylic acid.

acetic anhydride + salicylic acid → aspirin + acetic acid:


"C_4H_6O_3 +C_7H_7O_3 = C_9H_8O_4 +C_2H_4O_2"


The percent yield is the ration of the actual mass to theoretical yield:

"\% = \frac{m}{m_{theor}}\times100 \% = \frac{2.4 g }{3.6 g } \times 100 = 67\%"